-6.(Appendix) The following information is for Kershaw Cornpany for last month Budgeted direct labor mix at budgeted prices for actual output produced 3,825 skilled hours at $16 per hour 1.275 unskilled hours at $12 per hour 5,100 total hours Actual results 4,000 skilled hours at $19 per hour 1.000 unskilled hours at $9 per hour 5,000 total hours The direct labor yield variance for both types of labor together is: a. $1,500 favorable. b. $1,000 unfavorable. c. $1,000 favorable. d. $500 favorable. 7.(Appendix) Using the information in question 6, the mix variance for skilled labor is: a. $1,200 favorable. b. $4,000 unfavorable. C. $2.800 unfavorable
7.(Appendix) Using the information in question 6, the mix variance for skilled labor is: a. $1,200 favorable. b. $4,000 unfavorable. c. $2,800 unfavorable d. s2.800 favorable.
Answer #1
6) The direct labor yield variance for both types of labor together is
Actual Total Quantity of All Inputs Used × Budgeted Input Mix × Budgeted Price - Flexible Budget:Budgeted Total Quantity of All Inputs Allowed For Actual Output × Budgeted Input Mi×    x Budgeted Price
Skilled =   5,000 × 3875/5100 × $16 = 60000 Skilled =   5,100 × 3875/5100 × $16 = 61200
Unskilled = 5,000 x 1275/5100 x $12 15000 Unskilled = 5,100 x 1275/5100 x $12 15300
Total 75000 Total 76500
Direct labour Yield Variance = 75000 - 76500 1500 Favourable
7)
Using the information in question 6, the mix variance for skilled labor is
Mix Variance
Actual Total Quantityof All Inputs Used× Actual Input Mix× Budgeted Price Actual Total Quantity of All Inputs Used × Budgeted Input Mix × Budgeted Price
Skilled = 5,000 × 4000/5000 × $16 = 64000 Skilled =   5,000 × 3875/5100 × $16 = 60000
Mix Variance for Skilled Labor = 64000 - 60000 4000 Unfavourable